I need help doing this Identify the rate, base, and amount in the following applications. Do not solve the applications at this point
1. Ms. Jordan has been given a loan of $2500 for 1 year. If the interest charged is $275, what is the interest rate on the loan?
2.A survey of 400 people found that 66 were left-handed. What percent of those surveyed were left-handed?

Answer by learning b(2)

275/2500=r/100 =275 times 100 =R times 2500
R=27500/2500 = 11%